Main Logbooks page LST commissioning Shift checklists Safe state checklists Maintenance activities LST2-4 Commissioning Old logbooks
GENERAL CAMERA DRIVE ENERGY AMC
  LST1 Drive ELOG, all entries  ELOG logo
ID Date Author Type Category Subject Status Text Attachments
  279   Mon Feb 10 14:57:56 2020 Ino + Gael + ThierryOtherHardwareAzimuth Locking System FixedLast week we put in sevice the new ALS
completly modified:

_ Main System with new reducer.
 14x 
  278   Thu Jan 16 22:06:39 2020 tions applied to solve this issue :Martin Will, Thierry Le FlourTelescope OperationsSoftwareSemi-Expert Drive GUI not connecting / MOS server not workingFixedIn the late afternoon shifters noticed
the semi-expert drive GUI was not updating,
several restarts did not help. Checking into
 IMG-20200116-WA0020.jpegIMG-20200116-WA0029.jpeg 
  277   Sat Dec 7 19:16:01 2019 Daniel Mazin, Koji NodaHardwareHardwareALS East: problem and preliminary solution On Dec 6, Dario, Frederic and Yuji showed
to Koji Noda and Daniel Mazin how to operate
the new ALS. The idea was to do an unlock
 6x 
  276   Wed Nov 27 21:27:29 2019 Mathieu David ArmandTelescope Operations CDM & SG images with alternative target     
  275   Tue Nov 26 21:37:03 2019 David Sanchez, Armand Fiasson, Mathieu De BonyTelescope Operations CDM bending model image 26-11-2019 The aim was to observe several stars with
different magnitude and check the sensitivity
of the CDM. We wanted to use a semi-automatic
  
  274   Mon Nov 25 22:53:22 2019 Mathieu David ArmandTelescope Operations CDM images 19/11/25 Mg=1.25

DenebDark 60x1s
RA-2

TestTracking 60x1s Displacement
from Deneb RA-2

Deneb 60x1s

DenebLed
  
  273   Sun Nov 24 19:33:56 2019 Mathieu David ArmandTelescope OperationsGeneralCDM images for bending model - CDM images at park out position.
Exposure 100ms.
OARL on.
  
  272   Thu Sep 19 19:10:39 2019 Martin, Ino, Thierry, Quentin, ThomasTracking Debugging Star images with various Az and Zd offsets We tracked stars at various zenith angles
and changed the Az and Zd offsets, taking
images with the PSF camera for each position.
  
  271   Mon Sep 16 08:59:41 2019 Drive Team- Ino&Thierry + Ievgen + Vitalii + Jean-LucPLC + Control Software debugging Test TCU<->Drive Thierry&Ino on site

Ievgen, Vitalli and Jean-Luc remotely.

TCU tests to perform the parkin-in
  
  270   Thu Sep 12 16:41:53 2019 Drive Team- Ino&ThierryPLC + Control Software debugging   Tuesday September 110th :


Tests park-in and park-out
to manage the different type of failures 
Tests
  
  269   Sat Jul 27 11:11:08 2019 Yusuke, Satoshi, Tomo, YoshikiOther GUI stacked due to a wrong shutdown procedure Drive GUI stacked due to a wrong shutdown
procedure (Jul. 27th 01-02UTC). This was
simply a human error.
  
  268   Tue Jul 23 13:50:43 2019 Yusuke, PatriciaProblem Fixed Reboot Flywheel PLC Regarding the problem on the Flywheel PLC
(see
here), we entered UC1 under the supervision of Daniel
  
  267   Mon Jul 22 23:07:20 2019 Yusuke, Daniel K., Mitsunari, Satoshi, Tomo, Juan Abel, YoshikiTelescope Operations Flywheel PLC down After finishing the normal start-up procedure,
we tried to park out by GUI, but got an error
(22:30 local time): "MSG_WARNINGPark_Out_TimeOut"
 IMG_0165.jpgIMG_0167.jpg 
  266   Fri Jul 19 23:06:13 2019 Mitsunari, Tomohiro, Satoshi, Moritz, YusukeTelescope OperationsIssuesTelescope stuck in strange state/lost control by Semi-Expert GUI Tonight, we tracked Vega to test (a correction
of) the pointing offset of the Drive.

Around 22:30UTC, we tried different
  
  265   Tue Jul 9 00:46:52 2019 Mitsunari, Yukiho, Noah, MoritzTelescope OperationsIssuesPark-in problems At parking in, the telescope got stuck
at 94.254°. We checked in the container
that fuse Q21 went off. We did not see any
  
  264   Sat Jul 6 23:17:54 2019 Moritz, DirkTelescope Operations Park-in problems When parking in the telescope, accidently
hit "Shutdown" button instead of
"Parkin" in Drive Semi-epert GUI.
  
  263   Sat Jul 6 22:47:38 2019 Dirk HoffmannTelescope OperationsGeneralCar flash from road above+ When checking the telescope visually outside,
while it was in the "park-out"
position, a car happened to pass by from
  
  262   Fri Jul 5 23:23:57 2019 Dirk, MoritzTelescope OperationsGeneralPark-in report On park-in command, telescope went to the
right position 95/-1, but did not recognise
that it was there.
  
  261   Fri Jul 5 08:58:42 2019 Dirk HoffmannProblem FixedSoftwareSmall problem with park-in When parking in yesterday, the telescope
drove to the nominal position (approx. 95°
zenith angle and -1° azimuth), but then
  
  260   Wed Jul 3 21:02:09 2019 Dirk, Taka, Mitsunari, Yukiho, NoahTelescope OperationsIssuesFeedback from shiftcrew (For Thierry with love <3 :-)

The semi-expert drive control panel
uses various terms for the zenith angle:
  
  259   Tue Jul 2 21:25:07 2019 Dirk HoffmannApplicationsSoftwareCurrent & Power Monitoring stuck ("Socomec display") The Current&Power Monitoring was not
updated and hidden by two emtpy modal dialogue
windows.
  
  258   Fri Jun 28 23:06:36 2019 Dirk HoffmannTelescope OperationsGeneralTelescope stuck Ino explained: MOS was totally desynchronised
with PLC. (It was assuming that the telescope
was parked.) He parked the telescope manually.
  
  257   Fri Jun 28 22:49:39 2019 Dirk HoffmannTelescope OperationsGeneralTelescope stuck After tracking Mk421, we abort current
action (acknowledged) and try to go to a
new source: RA=308.1, Dec=40.96. Drive Control
  
  256   Thu Jun 27 12:19:02 2019 InoProblem FixedGeneralPLC Status During this weeks we solve this problems:

_ Park In: Sometimes the final
locking position was too limit and the "In
  
  Draft   Thu Jun 27 12:04:43 2019 InoProblem FixedGeneral  During this 2 weeks we solve this problems:

_ Park In: Sometimes the final
locking position was too limit and the "In
  
  254   Wed Jun 26 20:06:22 2019 InoPLC + Control Software debugging Tracking Trajectory We sent a tracking with with RA211,08 and
DEC29,52

1st time the goto position of azimuth
  
  253   Thu May 30 13:18:04 2019 AlessandroTelescope Operations semi expert gui (marginal) issues - During shutter tests, we were requested
to move the telescope many times between
two positions at different zenith. Sometimes,
  
  252   Fri Apr 26 23:17:07 2019 Alessandro for the drive teamTelescope OperationsIssuesdrive tracking issue 
We
encountered a couple of issue with tracking

- When we tried to track on polaris
  
  251   Thu Apr 18 21:21:25 2019 Drive Team- Ino&Thierry&AlessandroTelescope OperationsHardwareElevation limit switch issue 

The 2 swicthes are too far
 ElevationEndSwicth.png 
  250   Thu Apr 18 21:07:20 2019 Drive Team- Ino&Thierry&AlessandroTelescope Operations TESTS Parkout Parkin for fixing the bolt issue 
Different ParkOut and Parkin done
with different values for testing the torque
limitation. 

By decreasing
  
  249   Thu Apr 18 09:10:29 2019 InoPLC Software and ConfigurationHardwareOffset of the Zero Azimuth After pointing Polaris yesterday, it seems
that the zero position of the telescope was
incorrect but It's never had been calibrated
  
  248   Wed Apr 17 21:26:13 2019 Drive Team- Ino&Thierry&Alessandro, SatoshiTelescope OperationsGeneralTracking for mirrors alignment 
Tracking at position :

RA
= 101.288541 RA Offset = -0.6
DEC
= -16.71343 DEC Offset = 0.7


  
  247   Tue Apr 16 14:34:22 2019 Drive Team- InoTelescope OperationsHardwareCabling the azimuth end switches 
Cabling the azimuth end switches.
  
  246   Tue Apr 16 10:10:55 2019 Drive Team- Ino&ThierryPLC Software and Configuration Elevation setup 
1 degree has been added to the
Elevation accordingly to the last measurement
given by Olger
  
  245   Mon Apr 1 07:43:23 2019 Drive Team- Ino&Thierry&ArmandProblem FixedHardwareLST Parking control 
On Wenesday morning (27/03/19)
:

With Holger and Robert (MPI)
, many telescope "park in/park out"
  
  244   Mon Apr 1 07:36:04 2019 Drive Team- Ino&Thierry&ArmandTelescope OperationsGeneralTracking debug and setup 
Regulation of the motors (PID)
for slow and fast tracking and motor oscillations.
Implementation
and Debugging of the Semi-Expert control
  
  243   Thu Mar 28 10:34:41 2019 Drive Team- Ino&Thierry&Armand + Juli (IFAE)Telescope Operations High speed drive tests new session of high speed drive tests after
issues with Power Supply.


Tests of emergency stop
  
  242   Fri Mar 8 09:25:22 2019 Drive Team- Ino&Thierry&ArmandPLC + Control Software debuggingSoftwareTracking debug and setup 
Tracking procedure has been debugged
from the GUI-MOS (OPCUA Server with tracking
plugin)-PLC
Still some improvments to
  
  241   Fri Mar 8 09:20:29 2019 Drive Team- Ino&Thierry&ArmandTelescope OperationsGeneralHigh speed drive tests 
On Thrusday :

new
movements at 45% from -90 to 270 have been
done.
No possible movement at
  
  240   Wed Mar 6 16:30:56 2019 Drive Team- Ino&Thierry&ArmandPLC + Control Software debuggingSoftwareTracking mode debug 
Debugging of the MOS plugin corresponding
to the tracking 
  
  239   Wed Mar 6 15:33:14 2019 Drive Team- Ino&Thierry&ArmandTelescope OperationsGeneralTest High Speed Drive 
Test at 60% maximal speed from
-90 to 0

 Check of the
FW discharge : No error (FW 2 : 95% FW1 :
  
  238   Wed Mar 6 14:58:51 2019 Drive Team- Ino&Thierry&ArmandTelescope OperationsGeneralAccess tower failure During a park-in action , the acces tower
did not close correctly.
  
  237   Wed Mar 6 10:54:33 2019 Drive Team- Ino&Thierry&ArmandTelescope OperationsGeneralDRIVE : High Speed tests 
The FW issue is not solved . No
movement can be done
The FW is
going to be restarted ==> It takes several
  
  236   Tue Mar 5 14:42:22 2019 Drive Team- Ino&ThierryTelescope OperationsGeneralHigh speed drive tests . Fisrt day of the high speed drive tests
with IFAE team.


Tests for tuning and generating
  
  235   Thu Feb 28 20:01:44 2019 Drive Team- Ino&ThierryTelescope OperationsGeneralPolaris point for mirrors aligment Same operations as yesterday ==> Polaris
position
  
  234   Thu Feb 28 19:56:55 2019 Drive Team- Ino&ThierryPLC Software and Configuration Tracking debug and setup Debugging of the tracking sequence :


PLC sequence : trajectory
tracking data correctly managed
PLC
  
  233   Wed Feb 27 20:33:17 2019 Drive Team- Ino&ThierryTelescope OperationsGeneralPolaris point for mirrors aligment Same operations as yesterday ==> pointing
to Polaris with the same coordinates
  
  232   Tue Feb 26 20:47:15 2019 Drive Team- Ino&ThierryTelescope OperationsGeneralPolaris point for mirrors aligment Pointing to polaris ==>

Azimuth = 0.808

Elevation = 59.941 + 1 degree (correction)
  
  231   Tue Feb 26 20:45:41 2019 Drive Team- Ino&ThierryTelescope OperationsGeneralTracking debug Start of the tracking debug ==> PLC
and MOS 
  
  230   Wed Jan 30 15:43:12 2019 Drive TeamTelescope OperationsSoftwareTests Tracking Tests and debugging of the tracking procedure.
Many components are involved : PLC MOS GUIs.

Each component needs to be well
  
  229   Mon Jan 28 15:27:35 2019 Drive TeamTelescope OperationsSoftwareError Tests  
Debugging of the "Parkout"
sequence with errors ==> 

MOS
received the error and set the FSM in error
  
  228   Sat Jan 26 09:49:34 2019 Drive TeamTelescope OperationsGeneralGoto Position tests Many differents succesfull  "goto
position" operations ... MOS configuration
file (XML) modified to aplly to correct Azimth
  
  227   Fri Jan 25 14:12:28 2019 Drive TeamTelescope OperationsGeneralDrive Movement : PArkin and Parkout Many parkin operations done. To confirm
the different movements , a Parkin test with
a azimuth position 20 and Elevation 70.
  
  226   Fri Jan 25 12:39:52 2019 Drive TeamProblem FixedHardwareBrakes issues fixed The brake of the elevation issue has been
fixed. Operations on drive can be resumed.
  
  225   Thu Jan 24 10:52:59 2019 Drive TeamHardwareIssues  After issues during the parkin procedures
( 100% elveation torque ) :


a visual inpsection has
  
  224   Wed Jan 23 15:10:34 2019 Drive TeamTelescope OperationsHardware  Emergency stop applied because of a azimuth
value(encoder) not in the pre-defined range
+-0.015
  
  223   Wed Jan 23 11:26:29 2019 Drive TeamPLC Software and ConfigurationGeneralAzimuth tuning  
To define precisely the correction
to apply to Azimuth positionning, a set of movements
have been done to calculte the correction
 description_drive_pointe.pdf 
  3   Tue Jan 22 15:55:25 2019 Drive TeamSoftware InstallationSoftwareTESTS Parkout Parkin 
Many parkin and parkout operations
with different torque  (with the clamping
system of the access tower (from 18% to 8%
  
  2   Tue Jan 22 09:37:03 2019 Ino MonteiroPLC Software and ConfigurationSoftwareInstallation new PLC version Installation of the new PLC code and re-initilalisation
of the system.
  
  1   Fri Jan 18 10:14:51 2019 Thierry Le FlourSoftware InstallationGeneralInitlisation of the DRIVE logbook Here is the first entry in the DRIVE ELOG
logbook
  
ELOG V3.1.5-fc6679b